OpenStudy (anonymous):

Two lines, A and B, are represented by the following equations: Line A: 3x + 3y = 12 Line B: x + y = 4 Which statement is true about the solution to the set of equations? It is (12, 4). There are infinitely many solutions. It is (4, 12). There is no solution.

OpenStudy (anonymous):

please help

OpenStudy (michele_laino):

Hint: If I divide both sides of the first equation, by 3, I get: \[\begin{gathered} \frac{{3x + 3y}}{3} = \frac{{12}}{3} \hfill \\ \hfill \\ x + y = 4 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

is is it B

OpenStudy (michele_laino):

namely: line A and line B are two identical lines

OpenStudy (michele_laino):

that's right! Your answer is correct!

OpenStudy (anonymous):

yay can you help me on another one?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Two lines, A and B, are represented by the following equations: Line A: y = x − 1 Line B: y = −3x + 11 Which of the following options shows the solution to the system of equations and explains why? (3, 2), because the point does not lie on any axis (3, 2), because one of the lines passes through this point (3, 2), because the point lies between the two axes (3, 2), because both lines pass through this point

OpenStudy (anonymous):

@Michele_Laino can you help

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

for example, if I set x= 3 into the first equation, I get: y=3-2=... what is y?

OpenStudy (michele_laino):

oops...I get: y=3-1=...

OpenStudy (michele_laino):

what is y?

OpenStudy (anonymous):

2?

OpenStudy (michele_laino):

ok! So point (3,2) belongs to the first line

OpenStudy (anonymous):

I think it is D

OpenStudy (michele_laino):

Now, if I set x=3, into the second equation, I get: y=-3*3+11=... what is y?

OpenStudy (michele_laino):

that's right! Your option is the correct one!

OpenStudy (anonymous):

yay last one

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

Line M is represented by the following equation: x − y = 8 Which equation completes the system that is satisfied by the solution (18, 10)? 2x − y = 26 x + y = 18 2x − 2y = 36 x − y = −28

OpenStudy (michele_laino):

for example, if I set x=18 and y=10 into the last equation, I get: left side = 18-10=8 right side = -28 so 8 = -28 which is a false statement

OpenStudy (anonymous):

Is it B then because 10+8=18

OpenStudy (michele_laino):

B, no since we have: left side = x+y=18+10= 28 and right side = 18

OpenStudy (anonymous):

then C

OpenStudy (anonymous):

or A

OpenStudy (michele_laino):

why A?

OpenStudy (michele_laino):

hint: left side =2x-y=2*18 -10 =... please complete

OpenStudy (michele_laino):

what is 2*18 -10=...?

OpenStudy (anonymous):

26

OpenStudy (anonymous):

Sorry this is really hard I don't know this but how you are explaining I am understanding

OpenStudy (anonymous):

You still there???

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

So it is C?

OpenStudy (anonymous):

A is correct And thanks got a hundred percent

OpenStudy (michele_laino):

no, it is A, since the left side is 26, as you have computed, and the right side is 26 too

OpenStudy (michele_laino):

thanks! :) congratulation for your ratings!

OpenStudy (anonymous):

np have a great day

OpenStudy (michele_laino):

:)